OpenStudy (ichigomewmew):

subtracting rational expressions with different denominators 3x-2y 3x-y ------ - -------- + 2 6x 4x

OpenStudy (ichigomewmew):

so far I have this but I believe its wrong because I am unsure of the signs 3x-2y 3x-y ------ - -------- + 2 6x 4x = 3x-2y (2) 3x-y (3) ------ - -------- + (2) (6x) 6x (2) 4x (3) ( 2) (6x) 6x-4y - 9x-3y + 12x ______ _______ ___ 12x 12x 12x 6x-4y - 9x-3y + 12x ---------------------- 12x

OpenStudy (a1234):

So there is a +2 at the end?

OpenStudy (a1234):

If so, you need to multiply 2/1 by 12x since that is what the other denominators turn out to be.

OpenStudy (mathmale):

What y ou (Ichigo) have done looks fine to me.

OpenStudy (retireed):

3x-2y 3x-y ------ - -------- + 2 6x 4x = 3x-2y (2) 3x-y (3) ------ - -------- + (2) (12x) 6x (2) 4x (3) ( 1) (12x) Corrected here 6x-4y - 9x-3y + 24x corrected here ______ _______ ___ 12x 12x 12x 6x-4y - 9x-3y + 24x (6x-9x+24x) + (-4y +3y) 21x - y ---------------------- = -------------------- = --------- 12x 12x 12x